wiki:UC20190204

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 2/04/19

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Misty Bentz (GSU), John Bally (CU), Chip Kobulnicky (UWy), Eric Bellm (UW), Ben Williams (UW), Kal Kadlec (UW), Joanne Hughes (Seattle U), Moire Prescott (NMSU), Bill Ketzeback (APO), Mary Beth Kaiser (JHU), John Wisniewski (OU), Anne Verbiscer (UVa), Leslie Hebb (NAPG), Russet McMillan (APO), Erica Ellingson (CU)


User feedback and comments from institutional representatives:

  • Wyoming - nothing to report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - nothing to report
  • Oklahoma - nothing to report, Don Terndrup from Ohio State is willing to help with KOSMOS commissioning
  • NMSU - nothing to report
  • NAPG - nothing to report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 2/04/19

0) Overview

December saw one of the largest single storm snowfalls in memory. The observatory survived well, however, there was a fair amount of time lost in December due to clouds, snow, and humidity. When skies permitted we were able to open. January wasn’t much of a now maker but included unusual winds, humidity and clouds for much of the month.

1) Telescope

No problems.

2) Instruments

DIS - We had a couple of vacuum servicing runs, the last of which included the installation of a getter material and refurbished vacuum valves. Upon recovery we noticed an increase in readout noise on the blue camera, which we believe could be a result of the getter addition. We reopened the DIS blue camera and found a cracked insulator at the cold head interface that likely was the source of the noise. However, upon cool-down the CCD is not responding, therefore we are starting by troubleshooting the readout electronics outside the dewar since those were power cycled. We also warmed and pumped on the red camera in situ. This seemed to improve the contamination issue with the red camera. We had a cryotiger compressor failure during cool down of the red camera.

TripleSpec is operational.

Agile is operational.

Echelle is operational. The new Finger Lakes slit viewer camera is in the process of being integrated into the ICC.

ARCTIC is operational.

NIC-FPS is operational. It has been giving us some minor readout issues that can be cleared with an init.

3) ARCSAT

Operational.

4) Site Issues

Visitors to the site are reminded to follow the directions to APO given on the observatory web site rather than relying on a GPS for directions.


Additional telescope and instrument discussion:

The winter weather at APO has been atypically poor this year, with one major snowstorm right before the holidays and otherwise little moisture but lots of clouds, humidity and wind. This has resulted in more lost hours due to weather than normal, but the telescope has generally been productive when open. As a reminder, winter travel to the mountain is frequently hazardous. Roads are sometimes closed (or should be), and even clear-looking roads can have icy patches on the north-facing slopes. If you have a class trip or group of visitors coming to the site, be sure to follow the driving directions on the APO web site (GPS directions can send you on dirt roads that are impassable in bad weather). When in doubt, call the site before making the trip to obtain an up-to-date weather forecast and driving conditions.

DIS was serviced in January, at which time several vacuum seals were replaced and a getter was installed. This did not initially yield much improvement on the red camera; it was subsequently repumped and reserviced, and early indications are that it is holding a vacuum better than it had been. The news regarding the DIS blue camera is less good; after servicing it came back with lots of readout noise. Bill discovered a crack in an insulating disk that provides electrical insulation between the cryotiger cold head and the CCD, so clearly the camera didn’t go back together the way it was after the servicing. The problem has since gotten substantially worse; the blue camera now has a constant bias level and it is not responding to any source of light. While this suggests that something happened with the detector inside the camera, we are first doing some troubleshooting outside of the dewar since they had been power cycled. During the course of this investigation Bill and Ed found some signal levels that were not correct, therefore some components need to be replaced on saddlebag electronics board. This is currently in the works but it is slow going due to the delicate nature of these tasks. We will know more by the end of this week and then will determine if the problem has been fixed or if we have to go into dewar. It is particularly impressive and appreciated that Bill and Ed are able to make some progress on the DIS electronics given that the instrument was not built here. Nancy agreed to send out an update as soon as anything else is known about the blue camera.


Q1 3.5-m scheduling:

Russet just updated the Q1 schedule. There is still some unassigned time in Q1 that people can request: OPEN time in Feb and Mar and some dark time associated with DD01.


Q1 ARCSAT scheduling:

There are still a couple of weeks that are unassigned in Q1. We just changed the week of Feb 18 from open to engineering because we are making progress on the control of guider but we need to make some physical changes to the setup in addition to performing some other engineering tasks as well. There is still one other week in late March that is currently unassigned. We may continue the engineering then if we can’t complete everything in February, but as of now users can still request that time. We had our first Target of Opportunity time with ARCSAT a couple of weeks ago, which worked out well despite the fact that we do not have a formal policy in place for how to handle ToO requests on ARCSAT.


Q2 3.5-m scheduling:

All institutional allocations for Q2 were sent out; the requests are due back by Feb. 26. We set aside some extra engineering time in Q2 for KOSMOS commissioning. Last year in Q2 we had lots of visiting groups with some difficulty coordinating the various groups. Thus if you plan to bring a class or group to the site let us know soon so we can begin to coordinate.


NA2 Baffle Project:

The UVa team had been working toward the goal of accommodating an ARCTIC field of view of 7.5 arcmin but John Wilson recently verified that the actual FOV of ARCTIC is 7.85’ . He is now making some minor adjustments/redesign of the current baffle. Once that is complete he will issue an RFP for the full stray light analysis by an optical engineering firm. Overall this is good progress, and Nancy expects that the redesigned NA2 baffle can be completed and fabricated in time for installation during the 2019 summer shutdown.


KOSMOS Update:

The UW team will soon schedule the KOSMOS Critical Design Review, likely for mid/late February. This is later than expected because they had less engineering help than anticipated. The electronics and cart are currently being designed; the design of the NA2 adapter is complete. The KOSMOS dewar is about a third of the way through being built; then it will be sent out for coating and then back in the lab for assembly.


Seeking Help with APO Wiki Pages

We recently noticed that the APO 3.5m users' wiki is lacking data reduction guides for NICFPS and Agile. Nancy asked for volunteers to help provide some of that information. Both Misty Bentz and John Bally volunteered to upload their NICFPS data reduction scripts (in IRAF) to the wiki. Agile usage has been minimal over the last year, so perhaps not unsurprisingly there were no volunteers to provide an Agile reduction guide. Ben volunteered to ask Anjum whether she had any documentation she could upload to the wiki.


Training and Education Working Group Update

Erica Ellingson, who chaired the TEWG, called into today's telecon to summarize the TEWG Recommendations document for the Users Committee reps. The TEWG had ~ 15 members with broad representation across ARC. They conducted a series of telecons over the summer (which was challenging due to limited availability) and worked through keeping careful meeting minutes and living documents that could be reviewed asynchronously. The motivation behind the resulting TEWG white paper is that many groups are coming to APO with different configurations, numbers of attendees, levels of preparation and expectations. The TEWG tried to organize some of the expectations and provide some recommendations for best practices. One overarching theme that pervaded nearly all of the TEWG discussions is that interaction between the APO staff and visitors is paramount, both ahead of time and during the trip. The APO staff are very willing to work with trip leaders to help provide the students the best experience possible. The document contains some recommendations specific to those bringing large groups of trainees to the site (e.g. the possibility of using off-site housing, where to gather the students during the actual observing, how to have people rotate through the various tasks of observing to keep everyone engaged, and how to handle groups who do NOT have the primary goal of being trained and certified for future solo observing).

Joanne stated that the TEWG turned out to be a very good way to share and learn from people with different perspectives (especially from PhD-granting institutions versus undergraduate-focused programs). Many people hadn’t considered the different ways that various departments want to use APO to teach their students. Russet said that she and Candace have been exploring ways in which the Observing Specialists can help more with on site groups, e.g. by scheduling an extra Obs Spec to be working when larger groups come to the site (this again emphasizes the need to coordinate early with the schedulers so that they can provide an extra obs-spec on site during class trips). The TEWG recommendations document is nearly final now and Nancy will post it on the users' wiki and possibly elsewhere on the APO web site as well.


Response to Call for White Papers for 2020 Decadal Survey

In light of the recent government shutdown the deadline for Astro2020 white papers has been postponed to March 11, 2019. Nancy has started the white paper on Overleaf, and will be planning a telecon with people who can contribute to flesh out the content a bit more. She plans to invite the following people to the telecon, although others are welcome to join too: Leslie Hebb, Misty Bentz, Eric Bellm, Zach Berta-Thomspon, John Keller, Sarah Tuttle, Jon Holtzman. Once the telecon is scheduled Nancy will announce it to 35m-general.


Discussion of 25th Anniversary / Strategic Planning meeting in 2019

Nancy is planning a Strategic Planning meeting in 2019 to discuss the future priorities for the 3.5m user community. This could be held in conjunction with a 25th anniversary celebration (the telescope came online in 1994!), or it doesn't have to be. Nancy proposed several options such as hold the meeting in conjunction with the summer AAS meeting (in St Louis in 2019), hold it at APO/NSO, hold it at NMSU, hold it at some other central location, etc. It will likely be in the summer. Nancy distributed a survey to seek input concerning the where and when; once the date and location are locked in, she will provide more specifics concerning the format of the meeting. Users Committee reps are asked to identify 1-2 people from their departments who would be likely to attend this meeting, and have them complete this survey: https://goo.gl/forms/n20R2ZHB6RkP3iyi1. The deadline for completing the survey is Feb. 18, 2019.


ACTION ITEMS:

Open action items from previous meetings:

  • Russet will update the calibration section of the ARCTIC users manual once the Q3/Q4 scheduling tasks are completed. STATUS: CLOSED. Russet did update it, in particular the section on darks because at least one version of the guide was based on the status of the instrument in its first year and the operating temperature has changed since then and that affected the darks. The updated sections include darks and flats, bias, overscan, and choice of readout modes. She is seeking feedback to make sure the added text is clear and concise; if she does not receive any feedback she will consider this done.
  • Nancy will send out a draft of a white paper for the 2020 Decadal for UC feedback prior to the next telecon. STATUS: OPEN.
  • Users Committee reps should share the development re: the ARCTIC shutter speed with their users and let us know if anyone notices any reduction in photometric accuracy. STATUS: OPEN. Joanne hasn’t had a photometric night so she is waiting for her next observing run (3 nights coming up soon) in order to examine this. John Wisniewski, Leslie Hebb, and their Penn State collaborators have been doing photometry with ARCTIC + the diffuser and they haven’t noticed any performance issues over the past month.
  • Users Committee reps are asked to spread the word that if anyone knows of previous K/COSMOS users at KPNO or Chile, please reach out to them to get involved as we get closer to commissioning KOSMOS on the 3.5m. STATUS: CLOSED. John Wisniewski reported that Don Terndrup at Ohio State is willing to help.
  • Nancy will disseminate portions of her slides from the ARC Board meeting. STATUS: OPEN.
  • Nancy will poll users on where/when to hold the 2019 strategic planning meeting. STATUS: CLOSED. See survey given above.
  • Nancy will poll Users Committee reps on what day/time to hold our telecons in the Spring 2019 semester. STATUS: CLOSED. Telecon schedule has been set through June 2019.

New action items from this meeting:

  • Nancy will send out an update as soon as anything else is known about the DIS blue camera. STATUS: OPEN.
  • Misty Bentz and John Bally will upload their NICFPS data reduction scripts to the wiki. STATUS: OPEN.
  • Ben Williams will ask Anjum whether she had any Agile data reduction documentation she could upload to the wiki. STATUS: OPEN.
  • Nancy will disseminate the TEWG final report. STATUS: OPEN.
  • Nancy will organize a telecon to discuss the Astro2020 white paper. STATUS: OPEN.
  • Users Committee reps should discuss the 2019 strategic planning meeting with their departments, identify possible attendees, and have them complete the survey before 2/18/19. STATUS: OPEN.

Next meeting:

The next Users Committee telecon will be on Tuesday March 5, 2019, at 10:00 am MST.

Note: See TracWiki for help on using the wiki.